# Title of the Paper

### v2

@@ 6.7 Claims in Example 2
Suppose $G$ is a core-periphery interaction network. Each agent $i$’s centrality is given by $c_{i}(\gamma,G)=1/(1-\gamma\beta)$ if $i$ is core and $c_{i}(\gamma,G)=(1+\gamma(\alpha-\beta))/(1-\gamma\beta)$ if $i$ is peripheral. The sum of all core agents’ in-degrees is $l\beta+m\alpha$, and the sum of all peripheral agents’ in-degrees is 00. Plugging these into the left-hand side of (18) yields
@@
This is immediate from the fact that $01-\beta(2\gamma\beta-1)>$ 0. ∎
@@
6.8 Proposition 4
To show this proposition, it suffices to focus on the three-player interaction network in Example 1 and show that for any $\gamma>1/\sqrt{2}$, there exist interaction parameters $(\alpha,\beta)\in[0,1)^{2}$ such that
$\frac{\partial f}{\partial\alpha}<0.$
Specifically, the partial derivative is computed as
$\frac{\partial f}{\partial\alpha}=2\gamma+2\gamma^{2}(\alpha-\beta)-2\gamma\beta+1=2\gamma^{2}\alpha-2\gamma(\gamma+1)\beta+2\gamma+1.$
The infimum of $\partial f/\partial\alpha$ with respect to $(\alpha,\beta)$ is then
$\inf_{(\alpha,\beta)\in[0,1)^{2}}\frac{\partial f}{\partial\alpha}=-2\gamma^{2}+1.$
Because $-2\gamma^{2}+1<0$ if and only if $\gamma>1/\sqrt{2}$, the result follows. ∎
6.9 Derivation of the unique linear equilibrium under $\mathcal{I}^{\dagger}$
Following the proof in Section 6.1, it is easy to show that the sum of equilibrium slopes must be one for each agent in any linear equilibrium under $\mathcal{I}^{\dagger}$. This observation immediately implies that agent $i\neq 1$’s unique linear equilibrium strategy must be $a^{\dagger}_{i}=x_{i}$. Plugging these into agent $1$’s first-order condition yields
@@ 6.9 Derivation of the unique linear equilibrium under $\mathcal{I}^{\dagger}$
6.10 Derivation of $U^{\dagger}_{i}(\tau_{x},\tau_{y},G)$ and $W^{\dagger}(\tau_{x},\tau_{y},G)$
@@ 6.11 Proposition 5
6.12 Derivation of compound inequality (25)
Suppose $G$ is a core-periphery interaction network. By substituting the simplified expressions for the left-hand side of (23) (derived in Section 6.3) and the right-hand side of (24) (derived in Section 6.7), the two conditions can be combined into the single compound inequality:

### v3

@@ A.7 Claims in Example 2
Suppose $G$ is a core-periphery interaction network. Each agent $i$’s centrality is given by $c_{i}(\gamma,G)=1/(1-\gamma\beta)$ if $i$ is core and $c_{i}(\gamma,G)=(1+\gamma(\alpha-\beta))/(1-\gamma\beta)$ if $i$ is peripheral. The sum of all core agents’ in-degrees is $l\beta+m\alpha$, and the sum of all peripheral agents’ in-degrees is 0. Plugging these into the left-hand side of (18) yields
@@
This is immediate from the fact that $1-\beta(2\gamma\beta-1)>0$. ∎
@@
A.8 Corollary 1
Suppose $n\geq 3$. I construct two interaction networks $G$ and $G^{\prime}$ such that $G\leq G^{\prime}$ and $\Delta W(\tau_{x},\tau_{y},G^{\prime})<0<\Delta W(\tau_{x},\tau_{y},G)$. First, when $G$ is the empty network, defined by setting $\alpha=\beta=0$ in the core-periphery structure of Example 2, $\Delta W(\tau_{x},\tau_{y},G)>0$ holds for any $\gamma<1$. Second, by the derivation in Example 2, the welfare effect is negative if the interaction parameters $(\alpha,\beta)$ belong to the set $\mathcal{H}(\gamma,l/m)$, defined by
$f(\alpha,\beta,\gamma)<-\frac{l}{m}(1-\beta(2\gamma\beta-1)).$
Consider the limit as $\gamma\to 1$ and $\beta\to 1$ with fixed $\alpha=1/2$. The left-hand side converges to $(1+(1/2-1))^{2}-1/2=-1/4$, while the right-hand side converges to 0. By continuity, there exists a sufficiently high $\gamma<1$ and parameters $(\alpha^{\prime},\beta^{\prime})\in(0,1)^{2}$ such that the inequality holds. Let $G^{\prime}$ be the network characterized by these parameters. Then, $\Delta W(\tau_{x},\tau_{y},G^{\prime})<0$. Since $G$ has zero weights and $G^{\prime}$ has strictly positive weights, $G\leq G^{\prime}$. Thus, increasing connectivity from $G$ to $G^{\prime}$ reverses the sign of the welfare effect. ∎
A.9 Derivation of the unique linear equilibrium under $\mathcal{I}^{\dagger}$
Following the proof in Section A.1, it is easy to show that the sum of equilibrium slopes must be one for each agent in any linear equilibrium under $\mathcal{I}^{\dagger}$. This observation immediately implies that agent $i\neq 1$’s unique linear equilibrium strategy must be $a^{\dagger}_{i}=x_{i}$. Plugging these into agent $1$’s first-order condition yields
@@ A.9 Derivation of the unique linear equilibrium under $\mathcal{I}^{\dagger}$
A.10 Derivation of $U^{\dagger}_{i}(\tau_{x},\tau_{y},G)$ and $W^{\dagger}(\tau_{x},\tau_{y},G)$
@@ A.11 Proposition 4
A.12 Derivation of compound inequality (27)
Suppose $G$ is a core-periphery interaction network. By substituting the simplified expressions for the left-hand side of (25) (derived in Section A.3) and the right-hand side of (26) (derived in Section A.7), the two conditions can be combined into the single compound inequality:
@@ A.12 Derivation of compound inequality (27)
A.13 Proposition 5
Differentiating $W(\tau_{x},\tau_{y},G)$ with respect to $\tau_{y}$ yields
$\frac{\partial W}{\partial\tau_{y}}=\tau_{x}^{-1}\left[\frac{\partial(1-\gamma)}{\partial\tau_{y}}S(\gamma)+(1-\gamma)\frac{dS(\gamma)}{d\gamma}\frac{\partial\gamma}{\partial\tau_{y}}\right].$
Using the definition $\gamma=\tau_{x}/(\tau_{x}+\tau_{y})$, the partial derivatives are calculated as
$\frac{\partial\gamma}{\partial\tau_{y}}$ $=-\frac{\tau_{x}}{(\tau_{x}+\tau_{y})^{2}}=-\frac{\gamma}{\tau_{x}+\tau_{y}};$
$\frac{\partial(1-\gamma)}{\partial\tau_{y}}$ $=\frac{\tau_{x}}{(\tau_{x}+\tau_{y})^{2}}=\frac{\gamma}{\tau_{x}+\tau_{y}}.$
Substituting these expressions into the derivative of welfare gives
$\frac{\partial W}{\partial\tau_{y}}$ $=\tau_{x}^{-1}\left[\frac{\gamma}{\tau_{x}+\tau_{y}}S(\gamma)-(1-\gamma)\frac{\gamma}{\tau_{x}+\tau_{y}}\frac{dS(\gamma)}{d\gamma}\right]$
$=\frac{\gamma}{\tau_{x}(\tau_{x}+\tau_{y})}\left[S(\gamma)-(1-\gamma)\frac{dS(\gamma)}{d\gamma}\right].$
Since $\gamma,\tau_{x},\tau_{y}>0$, $\partial W/\partial\tau_{y}<0$ if and only if
$S(\gamma)<(1-\gamma)\frac{dS(\gamma)}{d\gamma}.$ (39)
Next, I derive the expression for $dS(\gamma)/d\gamma$. Differentiating $S(\gamma)$ yields
$\frac{dS(\gamma)}{d\gamma}$ $=\sum_{i\in N}\left[\frac{\partial c_{i}}{\partial\gamma}\left((1-d_{i}^{in})c_{i}+2d_{i}^{in}\right)+c_{i}\left((1-d_{i}^{in})\frac{\partial c_{i}}{\partial\gamma}\right)\right]$
$=\sum_{i\in N}\frac{\partial c_{i}}{\partial\gamma}\left((1-d_{i}^{in})c_{i}+2d_{i}^{in}+(1-d_{i}^{in})c_{i}\right)$
$=2\sum_{i\in N}\frac{\partial c_{i}}{\partial\gamma}\left((1-d_{i}^{in})c_{i}+d_{i}^{in}\right).$
Using the identity $\partial\mathbf{c}/\partial\gamma=\gamma^{-1}(\mathbf{c}^{\mathbf{c}}-\mathbf{c})$, the scalar derivative for agent $i$ is $\partial c_{i}/\partial\gamma=\gamma^{-1}(c_{i}^{\mathbf{c}}-c_{i})$. Substituting this into the expression for $dS(\gamma)/d\gamma$:
$\frac{dS(\gamma)}{d\gamma}$ $=2\sum_{i\in N}\gamma^{-1}(c_{i}^{\mathbf{c}}-c_{i})\left((1-d_{i}^{in})c_{i}+d_{i}^{in}\right)$
$=S^{\prime}(\gamma),$
which completes the proof. ∎
A.14 Proposition 6
First, I derive the equilibrium condition. Differentiating the payoff function (32) with respect to $a_{i}$ and setting the expected marginal utility to zero yields
$(1-r_{i})(a_{i}-\mathbb{E}_{i}[\theta])-r_{i}\sum_{j\neq i}g_{ij}(a_{i}-\mathbb{E}_{i}[a_{j}])=0.$
Rearranging terms, I obtain
$\left(1-r_{i}+r_{i}\sum_{j\neq i}g_{ij}\right)a_{i}=(1-r_{i})\mathbb{E}_{i}[\theta]+\sum_{j\neq i}r_{i}g_{ij}\mathbb{E}_{i}[a_{j}].$
Let $K_{i}\equiv 1-r_{i}+r_{i}d_{i}^{out}$. Dividing both sides by $K_{i}$ leads to
$a_{i}=\frac{1-r_{i}}{K_{i}}\mathbb{E}_{i}[\theta]+\sum_{j\neq i}\frac{r_{i}g_{ij}}{K_{i}}\mathbb{E}_{i}[a_{j}].$
Defining $\tilde{g}_{ij}\equiv r_{i}g_{ij}/K_{i}$ and noting that $1-\sum_{j\neq i}\tilde{g}_{ij}=(K_{i}-r_{i}d_{i}^{out})/K_{i}=(1-r_{i})/K_{i}$, this equation simplifies to the first-order condition (33). Since this condition is identical to that of the baseline model (2) with $\tilde{G}$ replacing $G$, the unique linear equilibrium is given by $a_{i}^{*}=(1-\gamma)c_{i}(\gamma,\tilde{G})y+(1-(1-\gamma)c_{i}(\gamma,\tilde{G}))x_{i}$.
Next, following the derivation in the baseline model, expanding agent $i$’s payoff function (32) yields
$U_{i}(\tau_{x},\tau_{y},G)$ $=-(1-r_{i})\mathbb{E}[X_{i}^{2}]-r_{i}\sum_{j\neq i}g_{ij}\mathbb{E}[X_{i}^{2}-2X_{i}X_{j}+X_{j}^{2}]$
$=-K_{i}\mathbb{E}[X_{i}^{2}]-\sum_{j\neq i}r_{i}g_{ij}\mathbb{E}[X_{j}^{2}]+2K_{i}\sum_{j\neq i}\tilde{g}_{ij}\mathbb{E}[X_{i}X_{j}],$
where the second equality follows from the definitions of $K_{i}$ and $\tilde{g}_{ij}$. Subtracting the payoff in the absence of public information, $U_{i}(\tau_{x},0,G)=-\tau_{x}^{-1}(K_{i}+r_{i}d_{i}^{out})$, and substituting the variance and covariance equilibrium moments derived in (7), (8), and (9) yields the change in agent $i$’s payoff:
$\Delta U_{i}(\tau_{x},\tau_{y},G)$ $=\tau_{x}^{-1}(1-\gamma)\left(-K_{i}c_{i}(c_{i}-2)-\sum_{j\neq i}r_{i}g_{ij}c_{j}(c_{j}-2)+2K_{i}c_{i}(c_{i}-1)\right)$
$=\tau_{x}^{-1}(1-\gamma)\left(K_{i}c_{i}^{2}-\sum_{j\neq i}r_{i}g_{ij}c_{j}(c_{j}-2)\right),$
[…]
$\sum_{i\in N}\sum_{j\neq i}r_{i}g_{ij}c_{j}(c_{j}-2)$ $=\sum_{j\in N}c_{j}(c_{j}-2)\sum_{i\neq j}r_{i}g_{ij}$
$=\sum_{j\in N}d_{j}^{in,r}c_{j}(c_{j}-2),$
where $d_{j}^{in,r}=\sum_{i\neq j}r_{i}g_{ij}$. Thus, aggregate welfare decreases if and only if
$\sum_{i\in N}\left(K_{i}c_{i}^{2}-d_{i}^{in,r}c_{i}(c_{i}-2)\right)<0,$
as desired. ∎
A.15 Proof of Corollary 2
The welfare effect is positive if and only if $\sum_{i\in N}\Omega_{i}>0$, where $\Omega_{i}=c_{i}[(K_{i}-d_{i}^{in,r})c_{i}+2d_{i}^{in,r}]$. I show that this sum is positive in both cases.
Case 1: Homogeneous Intensity ($r_{i}=r$ and $d_{i}^{out}=1$).
In this case, $K_{i}=1$ and $d_{i}^{in,r}=rd_{i}^{in}$. The normalized matrix $\tilde{G}$ has elements $\tilde{g}_{ij}=rg_{ij}$, and its row sums are uniform as $\tilde{d}^{out}_{i}=r$. The centrality is also uniform across agents, given by $c_{i}(\gamma,\tilde{G})=c=(1-\gamma r)^{-1}$. Factoring out the constant $c$, I obtain
$\sum_{i\in N}\Omega_{i}=c^{2}\sum_{i\in N}(1-rd_{i}^{in})+2cr\sum_{i\in N}d_{i}^{in}.$
Since $\sum d_{i}^{in}=\sum d_{i}^{out}=n$, this simplifies to $nc^{2}(1-r)+2crn$, which is strictly positive for $r<1$.
Case 2: Homogeneous Direction ($g_{ij}=\frac{1}{n-1}$).
The total payoff weight in this case is $K_{i}=1-r_{i}+r_{i}=1$ since $d_{i}^{out}=1$ by assumption. The effective in-degree is the average intensity of neighbors:
$d_{i}^{in,r}=\sum_{j\neq i}r_{j}g_{ji}=\frac{1}{n-1}\sum_{j\neq i}r_{j}.$
It follows that $d_{i}^{in,r}<1$ for all $i$ since $r_{j}<1$ for all $j\neq i$ by assumption. The coefficient of the squared centrality term in $\Omega_{i}$ is $K_{i}-d_{i}^{in,r}=1-d_{i}^{in,r}>0$. Thus, every individual welfare contribution $\Omega_{i}$ is strictly positive, which completes the proof. ∎
A.16 Proposition 7
First, I derive the first-order condition for the efficient strategy. The planner maximizes aggregate welfare $w(a,\theta)\equiv\sum_{i\in N}u_{i}(a,\theta)$ subject to the constraint that each agent $i$’s action $a_{i}$ depends only on their local information set $\mathbf{s}_{i}$. Differentiating $w$ with respect to agent $i$’s strategy $a_{i}$ and setting the expected marginal welfare to zero yields
$\left(1-\sum_{j\neq i}g_{ij}\right)(a_{i}-\mathbb{E}_{i}[\theta])-\sum_{j\neq i}g_{ij}(a_{i}-\mathbb{E}_{i}[a_{j}])-\sum_{j\neq i}g_{ji}(a_{i}-\mathbb{E}_{i}[a_{j}])=0.$
The first two terms represent agent $i$’s private marginal utility, while the third term captures the externality $i$’s action imposes on neighbors $j$ who seek to coordinate with $i$. Rearranging terms, I obtain
$\left(1-\sum_{j\neq i}g_{ij}+\sum_{j\neq i}g_{ij}+\sum_{j\neq i}g_{ji}\right)a_{i}=\left(1-\sum_{j\neq i}g_{ij}\right)\mathbb{E}_{i}[\theta]+\sum_{j\neq i}(g_{ij}+g_{ji})\mathbb{E}_{i}[a_{j}].$
Simplifying the coefficient on the left-hand side to $1+d_{i}^{in}$ and dividing both sides by this term leads to
$a_{i}=\frac{1-d_{i}^{out}}{1+d_{i}^{in}}\mathbb{E}_{i}[\theta]+\sum_{j\neq i}\frac{g_{ij}+g_{ji}}{1+d_{i}^{in}}\mathbb{E}_{i}[a_{j}].$
Defining the fictitious interaction weights $g_{ij}^{F}\equiv(g_{ij}+g_{ji})/(1+d_{i}^{in})$ and noting that
$1-\sum_{j\neq i}g_{ij}^{F}=1-\frac{\sum_{j\neq i}g_{ij}+\sum_{j\neq i}g_{ji}}{1+d_{i}^{in}}=\frac{1+d_{i}^{in}-d_{i}^{out}-d_{i}^{in}}{1+d_{i}^{in}}=\frac{1-d_{i}^{out}}{1+d_{i}^{in}},$
this equation simplifies to the desired first-order condition (36).
To ensure the existence of a unique linear efficient strategy, I verify that the spectral radius of $G^{F}$ satisfies $\gamma\rho(G^{F})<1$. Since $G^{F}$ is a non-negative matrix, its spectral radius is bounded by its maximum row sum. Using the assumption $\sum_{j\neq i}g_{ij}<1$, the row sum for any agent $i$ is strictly less than one:
$\sum_{j\neq i}g_{ij}^{F}=\frac{d_{i}^{out}+d_{i}^{in}}{1+d_{i}^{in}}=1-\frac{1-d_{i}^{out}}{1+d_{i}^{in}}<1.$
Thus, $G^{F}$ is row-substochastic, ensuring that the inverse $(I-\gamma G^{F})^{-1}$ exists and is non-negative. Since the condition (36) is identical to the equilibrium best-response condition (2) with $G^{F}$ replacing $G$, the unique linear efficient strategy is given by the unique linear equilibrium of the game defined by $G^{F}$. Applying Proposition 1 with $G^{F}$, agent $i$’s efficient strategy is thus
$a_{i}^{F}=(1-\gamma)c_{i}(\gamma,G^{F})y+(1-(1-\gamma)c_{i}(\gamma,G^{F}))x_{i},$
where $c_{i}(\gamma,G^{F})$ is the Katz-Bonacich centrality in the fictitious network $G^{F}$. ∎